OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points (-2,-16),(0,-4), (4,-28) a.y=2x^2-2x-4 b. y=4x^2 + 2x +2 c. y= -2x^2 + 2x -4 d. y= 2x^2 - 2x + 4

OpenStudy (whpalmer4):

A parabola has an equation which can be written in the form \[y = ax^2+bx+c\]We have 3 known points, and 3 unknown variables. It must be our lucky day, because 3 points will allow us to find all 3 unknown variables! Our points are: \((-2,-16),\,(0,-4),\,(4,-28)\) We can substitute each point into the equation to find a different equation in terms of \(a,b,c\): \[-16 = a(-2)^2 + b(-2) + c\]\[-4 =a(0)^2 + b(0) + c\]\[-28=a(4)^2 + b(4) + c\] Simplify and solve that system of equations to get the coefficients for your parabola equation.